HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

A SPECTROSCOPIC THERMOMETER: INDIVIDUAL VIBRATIONAL BAND SPECTROSCOPY WITH OH IN THE ATMOSPHERE OF WASP-33B

\underline{SERGEI N. YURCHENKO} , \underline{J. TENNYSON} , S. W.M. WRIGHT AND I. WALDMANN, Department of Physics and Astronomy, University College London,

Gower Street, WC1E 6BT London, United Kingdom; M. S. K. NUGROHO, Astrobiology Center, NINS, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan; M.

BROGI, Dipartimento di Fisica, Università degli Studi di Torino, via Pietro Giuria 1, I-10125, Torino, Italy; N.P. GIBSON, School of Physics, Trinity College Dublin, The University of Dublin, Dublin 2, Ireland; E.J. W. DE MOOIJ, Astrophysics

Research Centre, Queen’s University Belfast, Belfast BT7 1NN, UK

Individual vibrational band spectroscopy presents an opportunity to examine exoplanet atmospheres in detail by distinguishing where the vibrational state populations of molecules differ from the current assumption of a Boltzmann distribution. Here, retrieving vibrational bands of OH in exoplanet atmospheres is explored using the hot Jupiter WASP-33b as an example.^{1} We simulate low-resolution spectroscopic data for observations with the JWST’s NIRSpec instrument and use high resolution observational data obtained from the Subaru InfraRed Doppler instrument (IRD). Vibrational band-specific OH cross section sets are constructed and used in retrievals on the (simulated) low and (real) high resolution data. Low resolution observations are simulated for two WASP-33b emission scenarios: under the assumption of local thermal equilibrium (LTE) and a toy non-LTE model for vibrational excitation of selected bands.^{2}

We show that mixing ratios for individual bands can be retrieved with sufficient precision to allow the vibrational population distributions of the forward models to be reconstructed. For high resolution, cross-correlation applications, we apply the individual vibrational band analysis to an IRD spectrum of WASP-33b,^{3} applying an ‘un-peeling’ technique. Individual detection significances for the two strongest bands are shown to be in line with Boltzmann distributed vibrational state populations consistent with the effective temperature of the WASP-33b atmosphere reported previously. We show the viability of this approach for analysing the individual vibrational state populations behind observed and simulated spectra including reconstructing state population distributions.

1S. O.M. Wright et al., Astron., In press (2023); arXiv:2305.11071

2Sam O M Wright, Ingo Waldmann, and Sergei N Yurchenko, Non-local thermal equilibrium spectra of atmospheric molecules for exoplanets, Mon. Notices Royal Astron. Soc., 512, 2911 (2022)

3S.K. Nugroho et al., Astrophys. J. Lett., 910, L9 (2021)

p-number: p207  Submitted on Thu, 15 Jun 2023 22:27:35 +0200